Grade 10Mechanics

an automoblie is lifted by a hydraulic jack that consists of 2 pistons the large piston is 1m in diametre and small piston i 10cm in diametre if ‘W’ is the weight of the car how mich smaller a force is needed on the smal piston to lift the car?

Aiden Mathew Denny
8 Years agoGrade 10

2 Answers

Venkat
8 Years ago
dd
Venkat
ApprovedApproved Tutor Answer8 Years ago
The pressure on both pistons are same. Hence,
W/A1=F/A2
F=W(A2/A1)
F=W(PI (10/2)2/PI (100/2)2)
F=W(25/2500)
F=W/100